# A three-body form factor at sub-leading power in the high-energy limit: planar contributions

### v1

@@ 1 Introduction
During the past years, studies of small-mass factorization at the next-to-leading power (NLP) in $\lambda$ have emerged in the literature, either within the framework of soft-collinear effective theory (SCET) Bauer:2000ew ; Bauer:2000yr ; Bauer:2001ct ; Bauer:2001yt ; Beneke:2002ph ; Beneke:2002ni ; Hill:2002vw or based on diagrammatic analysis. The analysis of sub-leading soft emissions has been carried out in Larkoski:2014bxa . The construction of an easy-to-use helicity operator basis has been studied in Feige:2017zci ; Moult:2015aoa ; Kolodrubetz:2016uim . Anomalous dimensions of sub-leading power $N$-jet operators have been calculated in Beneke:2017ztn ; Beneke:2018rbh . Based on power counting and region analysis, the NLP factorization formula in the small-mass limit has been proposed for the Yukawa theory Gervais:2017yxv and for quantum electrodynamics (QED) Laenen:2020nrt ; terHoeve:2023ehm . An important ingredient in the QED factorization formula, the NLP jet function, has recently been computed in vanBijleveld:2025ekz . However, the NLP factorization formula has not been generically proven.
@@ 2 Setup of the calculation
Note that an integral family may contain several topologies, where a topology is defined according to which $D_{i}$’s appear in the denominator of Eq. (4) (i.e., $a_{i}>0$).
@@
$\mathcal{I}_{i}(x,y,z)=\sum_{n_{1},n_{2},n_{3}}\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)\,\epsilon^{n_{1}}\,x^{n_{2}}\,\log^{n_{3}}(x)\,.$ (7)
The purpose of this work then reduces to the calculation of the coefficients as a function of $y$ and $z$. We will constrain the highest orders of $x$, i.e., the integers $n_{2}$, for the purpose of calculating the squared-amplitude $\mathcal{F}^{(2)}$ up to $x^{1}$. This corresponds to the next-to-leading power (NLP) in the high-energy expansion. Nevertheless, our method can be used to compute higher power terms as well.
@@ 3 Differential equations and independent coefficients
where $\mathcal{P}^{(t)}_{i,j}$ are relational functions of $\epsilon$, $x$, $y$ and $z$. The differential equations with respect to $x$ lead to relations among the coefficients. Therefore, we only need to compute a set of independent coefficients, which is conceptually similar to the master integrals. After determining the set of independent coefficients, we can employ their differential equations with respect to $y$ and $z$ to obtain their analytic expressions.
@@ 3.1 Determination of expansion orders
Let’s focus on the equation with $i=11$, which arises from the differential equation of $\mathcal{I}_{11}$ with respect to $x$. In this case, one has $j\in\{1,8,11\}$ on the right-hand side of Eq. (11). For the NLP accuracy of the squared amplitude, one requires both $\mathcal{I}_{11}$ and $\mathcal{I}_{1}$ to be expanded to order $x^{3}$, which means that we need to consider the equation with $n_{2}=3$. However, the minimal value of $l$ for $i=11$, $j=1$ is $l=-2$. Therefore, the coefficient with $n_{2}-1-l=4$ will appear on the right-hand side of the equation. For this reason, we need to expand $\mathcal{I}_{1}$ up to order $x^{4}$.
Considerations similar to the above also applied to the constrains on $n_{1}$, i.e., the expansion orders in $\epsilon$. We need to perform such analysis for each sector, and update the constraints until all equations are closed.
@@ 4 Solving the differential equations
After constructing the differential equations, we now want to find the analytic solutions for the MCs. Ideally, it should be possible to transform the differential equations to a canonical form where the connection matrices $\mathcal{A}^{(y)}_{I,J}(y,z)$ and $\mathcal{A}^{(z)}_{I,J}(y,z)$ take a strictly triangular form. In that case the solutions can be easily written in terms of iterated integrals. However, in practice we find that it is not necessary, as the connection matrices are already in a block-triangular form with at most $3\times 3$ blocks. This can be iteratively solved as we will show in the following.
@@ 4.1 Simple cases
where $\mathcal{G}_{y}$ and $\mathcal{G}_{z}$ come from the already-solved MCs. If the coefficients $\mathcal{A}_{y}$ and $\mathcal{A}_{z}$ are both zero, the above equations can be easily solved by direct integration. We now consider the case where $\mathcal{A}_{y}$ and $\mathcal{A}_{z}$ are nonzero.
For convenience, we introduce the following functions
@@
where the constants of integration can be arbitrarily chosen. We then have the following relation:
$\mathcal{C}(y,z)=\frac{1}{\mathcal{M}_{y}}\left[\mathcal{F}_{y}(y,z)+\mathcal{C}_{z}(z)\right]=\frac{1}{\mathcal{M}_{z}}\left[\mathcal{F}_{z}(y,z)+\mathcal{C}_{y}(y)\right],$ (17)
where $\mathcal{C}_{y}(y)$ is a function that only depends on $y$, and similarly for $\mathcal{C}_{z}(z)$. We can readily derive a relationship between $\mathcal{C}_{y}$ and $\mathcal{C}_{z}$:
$\mathcal{C}_{y}=\frac{\mathcal{M}_{z}}{\mathcal{M}_{y}}(\mathcal{F}_{y}+\mathcal{C}_{z})-\mathcal{F}_{z}\,.$ (18)
By taking partial derivative of both sides with respect to $z$, we have
$\frac{\partial\mathcal{C}_{z}}{\partial z}=-\frac{\mathcal{M}_{y}}{\mathcal{M}_{z}}\frac{\partial}{\partial z}\left(\frac{\mathcal{M}_{z}}{\mathcal{M}_{y}}\right)\mathcal{C}_{z}-\frac{\mathcal{M}_{y}}{\mathcal{M}_{z}}\frac{\partial}{\partial z}\left(\frac{\mathcal{M}_{z}}{\mathcal{M}_{y}}\mathcal{F}_{y}-\mathcal{F}_{z}\right),$ (19)
which allows us to determine the analytic expression of $\mathcal{C}_{z}$ (as well as $\mathcal{C}_{y}$).
@@ 4.2 $2\times 2$ blocks
$\partial_{t}\mathcal{C}^{\prime}=\mathcal{A}^{\prime}_{t}\mathcal{C}^{\prime}+\mathcal{G}^{\prime}_{t}\,,\quad\mathcal{G}^{\prime}_{t}=b_{1}\mathcal{G}_{1}^{(t)}+b_{2}\mathcal{G}_{2}^{(t)}\,,$ (23)
which takes the form discussed in Section 4.1 and can be solved using the methods there. After obtaining the solution of $\mathcal{C}^{\prime}$, we can use it to derive analytic expressions of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Taking $\mathcal{C}_{1}$ as an example, we have
@@
Since $\mathcal{C}^{\prime}$ is known, the above equation can again be solve using the method of Section 4.1. The expression of $\mathcal{C}_{2}$ can then be obtained from those of $\mathcal{C}^{\prime}$ and $\mathcal{C}_{1}$.
@@ 4.3 $3\times 3$ blocks
One can observe that the two unknowns $b_{2}/b_{1}$ and $b_{3}/b_{1}$ are still coupled by the two equations, and this system is challenging to solve by brute force. Fortunately, the second method introduced in the previous subsection still works, and offers an algorithmic way to tackle this kind of problems. We make polynomial ansatze for $b_{1}$, $b_{2}$ and $b_{3}$, and solve for the coefficients from Eq. (4.3). After that, we can solve for $\mathcal{C}^{\prime}$ using the method of Section 4.1. We are then left with a $2\times 2$ system of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, which can be solved following the strategy of Section 4.2.
@@
A particular solution can be found by using the ansatze with $n=2$, and is given by
@@ 4.4 Boundary conditions and final results
The minimal powers of $\epsilon$ and $x$, and the maximal powers of $\log(x)$ for a given pair $(n_{1},n_{2})$ can be read off from the above expansion. Note that $f_{i,n_{1},n_{2},n_{3}}$ is just the coefficient $\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)$ at a particular kinematic point, and the general solution of $\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)$ has already been obtained from the differential equations. Such a solution can be written in terms of generalized polylogarithms (GPLs) with the help of PolyLogTools Duhr:2019tlz , up to some unknown constants. These constants can then be reconstructed as transcendental numbers by the PSLQ algorithm, using the high-precision numbers (about 100 decimal digits) of $f_{i,n_{1},n_{2},n_{3}}$. The transcendental basis is generated from the elements $\{\pi,\ln(2),\zeta_{2},\zeta_{3},\zeta_{4},\mathrm{Li}_{4}(1/2),\zeta_{5}\}$ up to weight 5. Note that the form of the analytic expressions can be different in different kinematic regions. We concretely work in the Euclidean region where all planar MIs have no imaginary part. The expressions in other kinematic regions can be obtained via analytic continuation.
@@ 5 Summary and outlook
There are several obvious follow-ups to this work in sight. Our method can be readily applied to non-planar contributions to this form factor. The main obstacle in such a calculation lies in the IBP reduction. We have attempted with the state-of-art reduction tools such as Kira, NeatIBP Wu:2023upw ; Wu:2025aeg and Blade Guan:2024byi , but have not succeeded within a reasonable amount of time. It will be necessary to develop more efficient reduction methods for such cutting-edge calculations. Another obvious generalization is to consider the $Q\bar{Q}g$ form factor in QCD, where genuine three-parton correlations can occur. This involves more integral families, but the method established in this work can still be straightforwardly applied. Finally, the differential equations for the master coefficients are solved by brute-force in this work. It will be interesting to investigate whether these differential equations can be casted into a canonical form which can be solved directly as iterated integrals. We leave these investigations to future works.

### v2

@@ 1 Introduction
During the past years, studies of small-mass factorization at the next-to-leading power (NLP) in $\lambda$ have emerged in the literature Penin:2014msa ; Larkoski:2014bxa ; Moult:2015aoa ; Kolodrubetz:2016uim ; Feige:2017zci ; Liu:2017vkm ; Beneke:2017ztn ; Gervais:2017yxv ; Beneke:2018rbh ; Liu:2018czl ; Liu:2019oav ; Laenen:2020nrt ; Liu:2020eqe ; Liu:2020tzd ; Liu:2020wbn ; Liu:2022ajh ; terHoeve:2023ehm ; vanBijleveld:2025ekz , either within the framework of soft-collinear effective theory (SCET) Bauer:2000ew ; Bauer:2000yr ; Bauer:2001ct ; Bauer:2001yt ; Beneke:2002ph ; Beneke:2002ni ; Hill:2002vw or based on diagrammatic analysis. The analysis of sub-leading soft emissions has been carried out in Larkoski:2014bxa . The construction of an easy-to-use helicity operator basis has been studied in Moult:2015aoa ; Kolodrubetz:2016uim ; Feige:2017zci . Anomalous dimensions of sub-leading power $N$-jet operators have been calculated in Beneke:2017ztn ; Beneke:2018rbh . Based on power counting and region analysis, the NLP factorization formula in the small-mass limit has been proposed for the Yukawa theory Gervais:2017yxv and for quantum electrodynamics (QED) Laenen:2020nrt ; terHoeve:2023ehm . An important ingredient in the QED factorization formula, the NLP jet function, has recently been computed in vanBijleveld:2025ekz . However, the NLP factorization formula has not been generically proven.
@@ 2 Setup of the calculation
Note that an integral family may contain several topologies, where a topology is defined according to which $D_{i}$’s appear in the denominator of Eq. (4) (i.e., $a_{i}>0$). The topologies belonging to each of the 5 families are illustrated in Fig. 1 of Appendix A.
@@
$\mathcal{I}_{i}(x,y,z)=\sum_{n_{1}=-2}^{\infty}\sum_{n_{2}=0}^{\infty}\sum_{n_{3}=0}^{n_{3,\max}}\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)\,\epsilon^{n_{1}}\,x^{n_{2}}\,\log^{n_{3}}(x)\,.$ (7)
The purpose of this work then reduces to the calculation of the coefficients as a function of $y$ and $z$. Note that the two-loop integrals considered in this work have no soft-collinear overlapped divergences, and the minimal power of $\epsilon$ is $-2$. The integrals also have no power-like singularities in the limit $m\to 0$, and therefore the minimal power of $x$ is 00. The maximal power of $\log(x)$, $n_{3,\max}$, depends on the value of $n_{1}$ and $n_{2}$. In practice, we will truncate the series in $\epsilon$ and $x$ for the purpose of calculating the squared-amplitude $\mathcal{F}^{(2)}$ up to $x^{1}$ and $\epsilon^{0}$. This corresponds to the next-to-leading power (NLP) in the high-energy expansion. Nevertheless, our method can be used to compute higher power terms as well.
@@ 3 Differential equations and independent coefficients
where $\mathcal{P}^{(t)}_{i,j}$ are rational functions of $\epsilon$, $x$, $y$ and $z$. The differential equations with respect to $x$ lead to relations among the coefficients. Therefore, we only need to compute a set of independent coefficients, which is conceptually similar to the master integrals. After determining the set of independent coefficients, we can employ their differential equations with respect to $y$ and $z$ to obtain their analytic expressions.
@@ 3.1 Determination of expansion orders
Let’s focus on the equation with $i=11$, which arises from the differential equation of $\mathcal{I}_{11}$ with respect to $x$. In this case, one has $j\in\{1,8,11\}$ on the right-hand side of Eq. (11). For the NLP accuracy of the squared amplitude, one requires both $\mathcal{I}_{11}$ and $\mathcal{I}_{1}$ to be expanded to order $x^{3}$, since the minimal power of $x$ in $\mathcal{A}_{1}$ and $\mathcal{A}_{11}$ is $-2$, which means that we need to consider the equation with $n_{2}=3$. However, the minimal value of $l$ for $i=11$, $j=1$ is $l=-2$. Therefore, the coefficient with $n_{2}-1-l=4$ will appear on the right-hand side of the equation. For this reason, we need to expand $\mathcal{I}_{1}$ up to order $x^{4}$.
Considerations similar to the above also applied to the constrains on $n_{1}$, i.e., the expansion orders in $\epsilon$. We need to perform such analysis for each sector, and update the constraints until all equations are closed. As a final outcome of such analyses, we provide the number of MCs for each integral family in Table 1.
Integral families $F_{1}$ $F_{2}$ $F_{3}$ $F_{4}$ $F_{5}$
Number of MCs 247 359 383 158 199
Table 1: The number of MCs in each integral family.
@@ 4 Solving the differential equations
After constructing the differential equations, we now want to find the analytic solutions for the MCs. Ideally, the solution procedure can be greatly simplified if the connection matrices $\mathcal{A}^{(y)}_{I,J}(y,z)$ and $\mathcal{A}^{(z)}_{I,J}(y,z)$ take a strictly triangular form. In that case the solutions can be easily written in terms of iterated integrals. This is similar in spirit to the idea of canonical differential equations Henn:2013pwa . However, the existing tools Gituliar:2017vzm ; Prausa:2017ltv ; Meyer:2017joq ; Lee:2020zfb ; Dlapa:2020cwj for finding canonical differential equations do not easily adapt to the current case. Fortunately, the connection matrices are already in a block-triangular form with at most $3\times 3$ blocks. These can be iteratively converted to a strictly triangular form, as will be demonstrated in the following.
@@ 4.1 Simple cases
where $\mathcal{G}_{y}$ and $\mathcal{G}_{z}$ come from the already-solved MCs. If the coefficients $\mathcal{A}_{y}$ and $\mathcal{A}_{z}$ are both zero, the above equations can be easily solved by direct integration. This corresponds to a strictly triangular system discussed earlier. We now consider the case where $\mathcal{A}_{y}$ and $\mathcal{A}_{z}$ are nonzero.
For convenience, we introduce the following functions:
$\mathcal{M}_{y}(y,z)=\exp\left(-\int\mathcal{A}_{y}(y,z)\,\mathrm{d}y\right),\quad\mathcal{M}_{z}(y,z)=\exp\left(-\int\mathcal{A}_{z}(y,z)\,\mathrm{d}z\right).$ (15)
Multiplying the MC by the two functions, the differential equations Eq. (14) are transformed into:
$\frac{\partial}{\partial y}\left(\mathcal{M}_{y}\mathcal{C}(y,z)\right)=\mathcal{M}_{y}\mathcal{G}_{y}(y,z)\,,\quad\frac{\partial}{\partial z}\left(\mathcal{M}_{z}\mathcal{C}(y,z)\right)=\mathcal{M}_{z}\mathcal{G}_{z}(y,z)\,.$ (16)
One can see that only already-solved MCs appear on the right-hand side of the above two equations. We would like to find a transformation function $\mathcal{T}(y,z)$ such that
$\frac{\partial}{\partial y}\left(\mathcal{T}\mathcal{C}(y,z)\right)=\mathcal{T}\mathcal{G}_{y}(y,z)\,,\quad\frac{\partial}{\partial z}\left(\mathcal{T}\mathcal{C}(y,z)\right)=\mathcal{T}\mathcal{G}_{z}(y,z)\,.$ (17)
This is an example of a strictly triangular system of differential equations, where the derivatives of the transformed MCs only depend on already-solved ones. Comparing Eq. (17) with Eq. (16), one can see that
$\mathcal{T}(y,z)=\mathcal{M}_{y}(y,z)\,f_{z}(z)=\mathcal{M}_{z}(y,z)\,f_{y}(y)\,,$ (18)
where the function $f_{z}(z)$ satisfies the differential equation
@@
and similar for $f_{y}(y)$. The fact that the right-hand side of the above equation is independent of $y$ follows from the compatibility condition of Eq. (14):
$\frac{\partial}{\partial y}\frac{\partial}{\partial z}\ln\frac{\mathcal{M}_{z}}{\mathcal{M}_{y}}=\partial_{z}\mathcal{A}_{y}-\partial_{y}\mathcal{A}_{z}=0\,,$ (20)
which further follows from the linear-independence of $\mathcal{C}(y,z)$ with respect to $\mathcal{G}_{y}$, $\mathcal{G}_{z}$ and their derivatives. From the above analysis, we can write
@@
Building on this result, we get general solution to the differential equaiton (17):
[…]
where $\mathcal{C}_{y}(y)$ is a function that only depends on $y$, and similarly for $\mathcal{C}_{z}(z)$. These two functions are related by the second equal sign in the above formula, and are only fixed once the arbitrary constant terms of the indefinite integrals are chosen. Using the fact that $\mathcal{C}_{y}(y)$ is independent of $z$, we can derive the differential equation of $\mathcal{C}_{z}(z)$ with respect to $z$:
[…]
which allows us to determine the analytic expression of $\mathcal{C}_{z}$ (as well as $\mathcal{C}_{y}$) up to a constant of integration. This constant should be fixed by a boundary condition, to be discussed later.
The above procedure of transforming the differential equations to a strictly triangular form can be naturally extended to $2\times 2$ blocks and $3\times 3$ blocks, as will be discussed in the following subsections.
@@ 4.2 $2\times 2$ blocks
$\partial_{t}\mathcal{C}^{\prime}=\mathcal{A}^{\prime}_{t}\mathcal{C}^{\prime}+\mathcal{G}^{\prime}_{t}\,,\quad\mathcal{G}^{\prime}_{t}=b_{1}\mathcal{G}_{1}^{(t)}+b_{2}\mathcal{G}_{2}^{(t)}\,.$ (27)
The above equation resembles the one studied in Section 4.1. We can employ the method there to get rid of the term $\mathcal{A}^{\prime}_{t}\mathcal{C}^{\prime}$, and effectively transform the differential equation to a strictly triangular form. After obtaining the solution of $\mathcal{C}^{\prime}$, we can use it to derive analytic expressions of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$. Taking $\mathcal{C}_{1}$ as an example, we have
@@
Since $\mathcal{C}^{\prime}$ is known, the above equation can again be solved using the method of Section 4.1. The expression of $\mathcal{C}_{2}$ can then be obtained from those of $\mathcal{C}^{\prime}$ and $\mathcal{C}_{1}$.
@@ 4.3 $3\times 3$ blocks
One can observe that the two unknowns $b_{2}/b_{1}$ and $b_{3}/b_{1}$ are still coupled by the two equations, and this system is challenging to solve by brute force. Fortunately, the second method introduced in the previous subsection still works, and offers an algorithmic way to tackle this kind of problems. We make polynomial ansatz for $b_{1}$, $b_{2}$ and $b_{3}$, and solve for the coefficients from Eq. (4.3). After that, we can solve for $\mathcal{C}^{\prime}$ using the method of Section 4.1. We are then left with a $2\times 2$ system of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, which can be solved following the strategy of Section 4.2.
@@
A particular solution can be found by using the ansatz with $n=2$, and is given by
@@ 4.4 Boundary conditions and final results
The minimal powers of $\epsilon$ and $x$, and the maximal powers of $\log(x)$ for a given pair $(n_{1},n_{2})$ can be read off from the above expansion. Note that $f_{i,n_{1},n_{2},n_{3}}$ is just the coefficient $\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)$ at a particular kinematic point, and the general solution of $\mathcal{C}_{i,n_{1},n_{2},n_{3}}(y,z)$ has already been obtained from the differential equations. Such a solution can be written in terms of generalized polylogarithms (GPLs) with the help of PolyLogTools Duhr:2019tlz , up to some unknown constants. These constants can then be reconstructed as transcendental numbers by the PSLQ algorithm, using the high-precision numbers (about 100 decimal digits) of $f_{i,n_{1},n_{2},n_{3}}$. The transcendental basis is generated from the elements $\{\pi,\ln(2),\zeta_{2},\zeta_{3},\zeta_{4},\mathrm{Li}_{4}(1/2),\zeta_{5}\}$ up to weight 5. It is worth mentioning that the DEs may demand higher order coefficients in $\epsilon$, as discussed in Sec. 3.1. That’s the reason why weight-5 functions and constants appear, despite the fact that two-loop amplitudes up to order $\epsilon^{0}$ only require weight-4 at most. Note that the form of the analytic expressions can be different in different kinematic regions. We concretely work in the Euclidean region where all planar MIs have no imaginary part. The expressions in other kinematic regions can be obtained via analytic continuation. Since our results are written in terms of GPLs, their analytic structure is very well understood. The branch cuts of GPLs are uniquely fixed by their symbols, and whenever one analytic continues across a branch cut, an extra term proportional to ±$\pm i\pi$ is generated, where the sign of $i\pi$ depends on the direction of going across the cut. See, e.g., Goncharov:2001iea ; Vollinga:2004sn ; zbMATH05494656 for more detailed discussions.
@@ 5 Summary and outlook
There are several obvious follow-ups to this work in sight. A rigorous verification of factorization at NLP necessitates the inclusion of non-planar contributions to the amplitudes. Our method can be readily applied to non-planar contributions to this form factor. The main obstacle in such a calculation lies in the IBP reduction. We have attempted with the state-of-art reduction tools such as Kira, NeatIBP Wu:2023upw ; Wu:2025aeg and Blade Guan:2024byi , but have not succeeded within a reasonable amount of time. It will be necessary to develop more efficient reduction methods for such cutting-edge calculations. Another obvious generalization is to consider the $Q\bar{Q}g$ form factor in QCD, where genuine three-parton correlations can occur. This involves more integral families, but the method established in this work can still be straightforwardly applied. Finally, the differential equations for the master coefficients are solved by brute-force in this work. It will be interesting to investigate whether these differential equations can be casted into a canonical form which can be solved directly as iterated integrals. We leave these investigations to future works.
@@ Appendix A Details of the integral families
(a) $F_{1}$ topology 1.
(b) $F_{1}$ topology 2.
(c) $F_{1}$ topology 3.
(d) $F_{1}$ topology 4.
(e) $F_{2}$ topology 1.
(f) $F_{2}$ topology 2.
(g) $F_{2}$ topology 3.
(h) $F_{2}$ topology 4.
(i) $F_{2}$ topology 5.
(j) $F_{2}$ topology 6.
(k) $F_{2}$ topology 7.
(l) $F_{2}$ topology 8.
(m) $F_{3}$ topology 1.
(n) $F_{3}$ topology 2.
(o) $F_{3}$ topology 3.
(p) $F_{3}$ topology 4.
(q) $F_{3}$ topology 5.
(r) $F_{4}$ topology 1.
(s) $F_{5}$ topology 1.